OpenStudy (anonymous):

3x – 2y = 7 9x – 6y = 3 What is the solution to the system?

OpenStudy (anonymous):

are you familiar with the elimination method ?

OpenStudy (anonymous):

the elimination method is when you eliminate a variable, either x or y, to solve for the other variable. To do this, there are times when you have to multiply an equation by a certain number so that the variable you are trying to eliminate, cancels out. So, which variable do you want to eliminate ? You pick.

OpenStudy (anonymous):

x

OpenStudy (anonymous):

now we will multiply the first equation by a -3. -3(3x - 2y = 7) this gets us : -9x + 6y = -21 are you with me so far ?

OpenStudy (anonymous):

so we have : -9x + 6y = -21 9x - 6y = 3 ---------------add 0 + 0 = -18 0 = -18 since the end result does not equal, this means there is NO SOLUTION.